# A Hybrid Anyon-Otto thermal machine

### v1

@@ Abstract
We propose a four-stroke quantum thermal machine based on the 1D anyon Hubbard model, which is capable of extracting the excess energy arising from anyon exclusion statistics at low temperature into finite work. Defining a hybrid anyon-Otto (HAO) cycle, we find that the low temperature work, in the absence of any interactions, is maximised in the pseudo fermionic limit, where the anyons most closely resemble free fermions. However, when weak interactions are introduced, the work output is no longer maximized at the bosonic or pseudo-fermionic extremes, but instead peaks at intermediate statistical angles. This clearly demonstrates that interactions and anyonic statistics conspire non-trivially to enhance performance, with interacting anyons offering greater quantum thermodynamic advantage than either bosons or pseudo-fermions, in this regime. Furthermore, we also identify different modes of operation of the HAO cycle, one of which emerges as a direct consequence of the finite anyon energy at low temperature.
@@
In recent times, a plethora of models of quantum thermal machines (QTMs) have been proposed and analyzed to understand the emergence of thermodynamic principles from quantum dynamics [1, 2]. The existence, albeit mostly theoretical, of such a large number of models can be attributed to two reasons. Firstly, the nondeterministic and destructive nature of generic quantum measurements make it difficult to uniquely and unambiguously define the quantum equivalents of classical work and heat. Secondly, genuine quantum phenomena have been shown to alter the performance of QTMs in a myriad of ways, depending on how they are incorporated into the model, for example, in the form of coherence in the system [3, 4], engineered heat baths [5, 6, 7, 8, 9, 10, 11], many-body effects [12, 13, 14, 15, 16, 17, 18, 19, 20], etc. Apart from theoretical investigations, progress has also been made on the experimental realization of QTMs [21, 22, 23, 24, 25].
@@
Figure 1: Schematic representation of the HAO cycle. The two unitary work strokes - expansion and compression are implemented through an explicit change of the Hamiltonian parameter $\lambda\in\{J,U\}$. The heat strokes consist of thermalization with one of the heat baths with temperature $T_{A}$ or $T_{B}$ and a change in the statistical parameter $\theta$.
@@
The simplest and most commonly studied models of QTMs are those based on the four-stroke quantum Otto cycle [2]. The standard Otto cycle consists of two work strokes and two thermalization strokes. In the former, a system is made to evolve unitarily through a quench in some parameter of the Hamiltonian, with the energy change of the system associated with work. The work strokes are interceded by the thermalization strokes in which the system evolves dissipatively in contact with a heat bath, the energy change being associated with the heat transferred. At the end of the full cycle, the system returns to its initial state with a net conversion of quantum heat into quantum work and vice versa. The direction of heat flow and work output determines the operating mode of the cycle, and the thermodynamically allowed modes are the engine, refrigerator, accelerator and heater. The engine (regfrigerator) mode is characterized by a net heat transfer from the hotter (colder) to the colder (hotter) bath with a net work output (input). In accelerator mode, input work is used to boost the heat transfer from the hotter to the colder bath, while in heater mode, input work is fully converted into heat and dumped into each of the baths.
@@
Figure 2: Work output $W$ ((a) and (d)), and heat exchanged with the baths $Q_{A}$ ((b) and (e)) and $Q_{B}$ ((c) and (f)), in the HAO cycle as a function of the bath temperatures $T_{A}$ and $T_{B}$. The top panel corresponds to the standard Otto cycle with $\theta_{1}=\theta_{2}=0$, showing the three possible working modes - engine(E), refrigerator (R) and accelerator (A). The bottom panel shows the emergence of the inverse accelerator mode (IA) in the low temperature limit for $\theta_{2}=0,\theta_{1}\neq 0$. The plots depicted above are obtained numerically for a chain of length $L=8$, with number of particles $N=4$, and $U=0$, $J_{1}=1.0$ and $J_{2}=2.0$.
@@
The Otto cycle in the engine mode, by definition, relies on thermal energy for work output; the latter thus vanishes in the limit of zero temperature of the heat baths. However, quantum statistical properties depend crucially on the nature of the particles at low temperatures and this has been recently exploited to design and experimentally realize the Pauli engine [23]. In this variant of the regular Otto engine, the thermalization strokes are replaced by the so called Pauli strokes, in which the system undergoes a bosonization or fermionization process. Specifically, rather than thermalizing with a heat bath, the statistical nature of the system is altered by tuning it from a molecular Bose condensate to a Fermi gas or vice-versa. By virtue of the Pauli exclusion principle, the system has a higher energy ground state in the fermionic state as compared to the bosonic state. The Pauli engine operates by extracting this difference in ground state energy, defined as the Pauli energy, into useful work. Importantly, the Pauli engine has no classical analogue and is purely driven by quantum statistical phenomena. To this end, we note that the relation between the performance of QTMs and statistical properties of the system has been explored in a number of other works, especially with respect to fermionic and bosonic statistics [26, 27, 28, 29, 30].
@@
It has been shown in recent times that quantum particles such as anyons [31, 32, 33, 34, 35, 36, 37, 38], including impurity excitations in fractional quantum Hall liquids which exhibit fractional angular momentum and effective anyonic statistics [39, 40], and paraparticles [41, 42], satisfy statistical properties different from bosons or fermions. The performance of QTMs based on systems satisfying such non-trivial statistics has largely remained unexplored, except recent models based on few-body anyonic systems [43, 44, 45, 46]. In this work, we define and analyze a hybrid Anyon-Otto (HAO) cycle (see Fig.1), based on the anyon Hubbard model (AHM) [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61], in which the statistical properties of the particles can be continuously tuned from boson-like to pseudo-fermion-like (satisfy anti-commutation relations only when particles on different sites). We contrast the operation of the HAO cycle with the Otto cycle in the same temperature and parameter regime except that the statistical properties of the system remain unchanged throughout the Otto cycle. In the limit of vanishing temperatures of the baths, we find that the HAO cycle can produce a finite work output due to the presence of an anyon energy, defined analogously to the Pauli energy. In the absence of explicit interactions between the anyons, we find that the HAO cycle can operate in an inverse accelerator mode for finite but small temperatures. This mode, characterized by a heat transfer from colder to hotter bath and a net work output, exhibits an apparent violation of the second law which stems from ignoring the work production in the anyonization stroke. When the anyons are made to interact weakly, we observe that anyonic statistics can enhance the low temperature work output as compared to bosonic or pseudo-fermionic statistics. Finally, we note that the anyon Hubbard model has already been experimentally realized recently [47, 57, 58], thus rendering the possibility of an experimental verification of the results discussed in the rest of the article.
@@
The 1D AHM is realized by introducing a synthetic gauge field in the hopping term of the Bose Hubbard model,
$H=-J\sum_{j}\left(b_{j}^{\dagger}e^{-in_{j}\theta}b_{j+1}+h.c.\right)+\frac{U}{2}\sum_{j}n_{j}\left(n_{j}-1\right),$ (1)
where $b_{j}$ are the bosonic annihilation operators, $n_{j}=b_{j}^{\dagger}b_{j}$ is the number operator, $J$ is the tunneling amplitude, $U$ is the interaction strength, and $\theta$ is a phase parameter. Defining the annihilation operators $a_{j}=e^{-i\theta\sum_{1\leq l\leq j-1}n_{j}}b_{j}$, the equivalent anyon model is obtained,
$H=-J\sum_{j}\left(a_{j}^{\dagger}a_{j+1}+h.c.\right)+\frac{U}{2}\sum_{j}n_{j}\left(n_{j}-1\right),$ (2)
@@
The HAO cycle that we propose here is motivated by the Pauli and Otto cycles. A hallmark of this cycle is the ability to function both as an anyon engine (defined analogously to the Pauli engine) at low temperature and as the standard Otto engine at high temperature. We recall that in the Otto cycle, the following definitions of work and heat are used, $W=\int{\rm Tr}[\rho\frac{\partial H}{\partial t}]dt$ and $Q=\int{\rm Tr}[\frac{\partial\rho}{\partial t}H]dt$, where $\rho$ is the density operator representing the state of the system. Thus, a unitary evolution with a time-dependent Hamiltonian only contributes to the work, while a dissipative evolution, such as thermalization with a heat bath, only results in heat transfer. We now define the HAO cycle as follows. At the start of the cycle, we assume that the system with Hamiltonian $H(\theta_{1},\lambda_{1})$ is in equilibrium with a thermal bath with temperature $T_{A}$. The initial state is therefore a thermal Gibbs state $\rho_{1}\equiv e^{-\beta_{A}H(\theta_{1},\lambda_{1})}/{\rm Tr}\left[{e^{-\beta_{A}H(\theta_{1},\lambda_{1})}}\right]$. The HAO cycle consists of the following strokes (see Fig. 1):
@@
    Thermalization B: ($\theta_{1}\to\theta_{2}$, contact with heat bath) - This stroke consists of two sub-strokes – the phase parameter is tuned from $\theta_{1}$ to $\theta_{2}$ followed by thermalization with a heat bath of temperature $T_{B}$. Note that the two sub-strokes can be carried out simultaneously provided that the phase parameter is changed over a time interval much shorter than the thermalization time-scale. The steady state reached is thus given by, $\rho_{3}=e^{-\beta_{B}H(\theta_{2},\lambda_{2})}/{\rm Tr}\left[{e^{-\beta_{B}H(\theta_{2},\lambda_{2})}}\right]$. The change in the energy expectation value in this stroke has two contributions, one from the energy dissipated to the heat bath and the other from the work done in changing the phase parameter. However, for reasons that we shall clarify below, we associate the total energy change in this stroke with heat transfer,
@@
    Thermalization A: ($\theta_{2}\to\theta_{1}$, contact with heat bath) - In the final stroke, the phase parameter is restored to its initial value $\theta_{1}$ and the subsequent thermalization with heat bath at temperature $T_{A}$ restores the system back to its initial state $\rho_{1}$. Thus,
@@
The HAO cycle defined above differs from the standard quantum Otto cycle in the two thermalization strokes. In the latter case, the statistical properties of the quantum model remain unchanged, and thus the heat change is solely associated with the energy gained or lost to the heat baths. In contrast, thermalization in the HAO cycle is accompanied by a explicit change of statistical properties through the parameter $\theta$. It is important to note that ramping the phase parameter modifies the Hamiltonian itself and thus technically amounts to performing a certain amount of work. However, in the anyon equivalent of the model defined in Eq. (2), $\theta$ implicitly controls the statistical properties and hence the equilibrium configuration of the system. Since the purpose of the HAO engine is to convert both thermal energy and anyon energy into useful work, we associate the total energy change in the modified thermalization strokes only with heat energy.
@@
We first consider the operation of the HAO cycle in the non-interacting limit $U=0$ with $\lambda\equiv J$. In this limit, the work strokes are always adiabatic and thus the cycle operation is independent of the stroke duration $\tau$. We also assume that the total number of particles remain conserved throughout the cycle. In Fig. 2-2, we examine the total work $W=W_{12}+W_{21}$, $Q_{A}$ and $Q_{B}$ as a function of bath temperatures $T_{A}$ and $T_{B}$ for the standard Otto cycle ($\theta_{1}=\theta_{2}=0$) with $J_{1}=2.0$, $J_{2}=1.0$, $L=8$ and $N=4$. Depending on the sign of these quantities and the relative magnitude of $T_{A}$ and $T_{B}$, three modes of cycle operation can be identified, namely the engine mode (E) with $Q_{A},W>0$, $Q_{B}<0$ for $T_{A}>T_{B}$, the refrigerator mode (R) with $Q_{A},W<0$, $Q_{B}>0$ for $T_{A}>T_{B}$ and the accelerator mode (A) with $Q_{A},W>0$, $Q_{B}<0$ for $T_{A}<T_{B}$. The E-R transition can be identified by the line $W=Q_{A}=Q_{B}=0$. For $T_{A},T_{B}\to 0$, the work output and the heat vanish identically.
Having established the modes of operation in the Otto cycle, we now consider the situation $\theta_{1}=\pi$, $\theta_{2}=0$, that is, the statistical properties of the system are altered between bosonic and pseudo-fermionic during the thermalization strokes. From Fig. 2-2, it is evident that in the large temperature limit of the baths, $T_{A},T_{B}\gg 1$, the HAO cycle is identical to the Otto cycle owing to the fact that the anyonic statistics can be well-approximated by Boltzmann statistics in this limit. However, at low temperatures, $T_{A},T_{B}\to 0$, the excess energy resulting from the development of anyonic exclusions begins to manifest itself in the form of the non-zero work output and heat exchanges. This leads to the emergence of an inverse accelerator mode - heat flows from the colder bath to hotter bath with a net work extraction, $Q_{A},W>0$, $Q_{B}<0$ for $T_{A}<T_{B}$.
The apparent violation of the second law in the thermodynamics in the inverse accelerator mode is due to the fact the work produced during the bosonization or anyonization strokes have been defined to be part of the heat source. In Fig. 3, we observe that the average work output per particle $W/N$ increases monotonically with increasing $\theta_{1}$ in the inverse accelerator mode ($T_{A}=0.2$, $T_{B}=0.4$), provided $T_{A}\to 0$. Similarly, the work input decreases with increasing $\theta_{1}$ in the accelerator mode ($T_{A}=0.2$, $T_{B}=0.8$) in the same limit of $T_{A}$. Note that it is sufficient for only $T_{A}$ to be small as the anyonization process occurs in contact with the heat bath of temperature $T_{A}$.
@@
We now consider the situation in which work is performed by ramping the interaction parameter $U$ during the work strokes, holding $J$ to a constant value. It is important to note that the Hamiltonian does not commute with itself at different times during the work strokes, which leads to non-adiabatic excitations for finite stroke duration $\tau$. However, for simplicity, we shall assume that $\tau\to\infty$ as the results discussed in the following do not change qualitatively in the presence of non-adiabatic excitations. In order to ensure that the anyon energy is not affected by the interaction, we choose $U_{1}=0$. The work output $W$ at low temperature, $T_{A}=T_{B}=0.1$, as a function of the statistical parameter $\theta_{1}$ is shown in Fig. 4 for different interaction strengths $U_{2}$. In the weakly interacting limit $U\ll J$, we observe that $W$ is maximized for $\theta_{1}=\theta^{*}$ such that $0<\theta^{*}<\pi$, for $N\geq L/2$. However, as the interaction strength is increased, the maxima shifts towards the pseudo-fermionic limit $\theta^{*}=\pi$.
@@
The higher probabilities of double or greater occupancy in the bosonic and pseudo-fermionic limit can be understood by reverting back to the bosonic Hamiltonian defined in Eq. (1). Since the hopping ampitude acquires a density dependent phase for any $\theta\neq 0$, the system minimizes energy by preferring lesser contributions from configurations in which a given site is doubly or more occupied. For less than half-filling $N<L/2$, this is easily achieved as there are sufficient number of empty sites to satisfy lesser or equal to unit occupancy per site. Thus, the probabilty for higher occupancy decreases monotonically with increasing value of $\theta$. For $N\geq L/2$, the situation is not straightforward and the system is forced to choose configurations with greater than unit occupancy for some sites. However, as $\theta\to\pi$, the density dependence phase reduces to unity even for higher occupancy, and thus the probability of more than unit occupancy increases again. To support the above argument, we plot in Fig. 5 the probability $P(n_{j})$ that the $j^{th}$ site is occupied by $n_{j}$ particles for a couple of nearby sites $j=3,4$ in the system with $L=8$, $J=1.0$ and $U=0$. Although $P(n_{j}=0,1)>P(n_{j}=2,3,4)$ for both the sites, $E_{per}$ has no contributions from the zero or unit occupancies as the interaction energy vanishes identically for $\braket{\psi_{0}|n_{j}|\psi_{0}}=0,1$. It can be seen from Fig. 5 that the $P(n_{j}=2)$ decreases initially as $\theta$ is increased from zero, before increasing again on approaching $\theta=\pi$. Importantly, the increase in the value of $P(n_{4}=2)$ is accompanied by an increase in the value of $P(n_{3}=0)$ and a decrease in the value of $P(n_{3}=1)$, suggesting that double occupancy is preferred at the expense of unit occupancy in the limit $\theta\to\pi$.
Reconciliation with second law
In the previous sections, we have argued that the apparent violation of second law of thermodynamics in the inverse accelerator mode results from the identification of the anyon energy as a heat source. To support this argument further, we modify the definition of work in Eq. (3) and Eq. (5) to incorporate the energy change associated with the anyonization and bosonization processes. Likewise, we subtract the same from the heat exchanges to ensure consistency with the first law. Furthermore, in this case, we assume that the anyonization/bosonization process is carried out much faster than the time-scale of thermalization with the baths. This ensures that energy change during the anyonization/bosonization process is only due to work performed by explicit change of the parameter $\theta$ in the Hamiltonian. Thus, the redefined work and heat are, $\overline{W}_{12(21)}={W}_{12(21)}-\Delta_{B(A)}$, $\overline{W}=\overline{W}_{12}+\overline{W}_{21}$, $\overline{Q}_{A(B)}=Q_{A(B)}-\Delta_{A(B)}$, where

### v2

@@ Abstract
We propose a four-stroke quantum thermal machine based on the 1D anyon Hubbard model, which is capable of extracting the excess energy arising from anyon exclusion statistics at low temperature into finite work. Defining a hybrid anyon-Otto (HAO) cycle, we find that the low-temperature work, in the absence of any interactions, is maximized in the pseudo-fermionic limit, where the anyons most closely resemble free fermions. However, when weak interactions are introduced, the work output is no longer maximized at the bosonic or pseudo-fermionic extremes but instead peaks at intermediate statistical angles. This clearly demonstrates that interactions and anyonic statistics conspire non-trivially to enhance performance, with interacting anyons offering greater quantum thermodynamic advantage than either bosons or pseudo-fermions, in this regime. Furthermore, we also outline an experimental protocol to realize the HAO cycle using ultracold atoms in an optical lattice.
I Introduction
@@
In recent times, a plethora of models of quantum thermal machines (QTMs) have been proposed and analyzed to understand the emergence of thermodynamic principles from quantum dynamics [kosloff14, souravReview21]. The existence, albeit mostly theoretical, of such a large number of models can be attributed to two reasons. Firstly, the nondeterministic and destructive nature of generic quantum measurements makes it difficult to uniquely and unambiguously define the quantum equivalents of classical work and heat. Secondly, genuine quantum phenomena have been shown to alter the performance of QTMs in a myriad of ways, depending on how they are incorporated into the model, for example, in the form of coherence in the system [hardal15, camati19, mohan25], engineered heat baths [scully03, lutz14, manzano16, klaers17, Niedenzu18, ray23, bera24, sen25], many-body effects [campisi16critical, fazio16, hartmann20, revathy20, piccitto22, solfanelli23, arezzo24, mukherjee24, brollo25], etc. In addition to theoretical investigations, progress has also been made in the experimental realization of QTMs [von19, gluza21, Koch2023, Aamir2025, uusnäkki25experiment].
@@
Figure 1: Schematic representation of the HAO cycle. The two unitary work strokes - expansion and compression, are implemented through an explicit change of the Hamiltonian parameter $\lambda\in\{J,U\}$. The heat strokes consist of thermalization with one of the heat baths with temperature $T_{A}$ or $T_{B}$ and a change in the statistical parameter $\theta$.
@@
The simplest and most commonly studied models of QTMs are those based on the four-stroke quantum Otto cycle [souravReview21]. The standard Otto cycle consists of two work strokes and two thermalization strokes. In the former, a system is made to evolve unitarily through a quench in some parameter of the Hamiltonian, with the energy change of the system associated with work. The work strokes are interceded by the thermalization strokes in which the system evolves dissipatively in contact with a heat bath, the energy change being associated with the heat transferred. At the end of the full cycle, the system returns to its initial state with a net conversion of quantum heat into quantum work or vice versa. The direction of heat flow and work output determines the operating mode of the cycle, with the thermodynamically allowed (respecting the second law) modes being the engine, refrigerator, accelerator, and heater. The engine (refrigerator) mode is characterized by a net heat transfer from the hotter (colder) to the colder (hotter) bath, with a net work output (input). In accelerator mode, input work is used to boost the heat transfer from the hotter to the colder bath, while in heater mode, input work is fully converted into heat and dumped into each of the baths.
@@
Figure 2: Work output $W$ ((a) and (d)), heat exchanged with the baths $Q_{A}$ ((b) and (e)), and $Q_{B}$ ((c) and (f)), in the HAO cycle as a function of the bath temperatures $T_{A}$ and $T_{B}$. The top panel corresponds to the standard Otto cycle with $\theta_{1}=\theta_{2}=0$, showing the three possible working modes - engine(E), refrigerator (R) and accelerator (A). The bottom panel shows the emergence of the inverse accelerator mode (IA) in the low temperature limit for $\theta_{2}=0,\theta_{1}\neq 0$. The plots depicted above are obtained numerically for a chain of length $L=8$, with number of particles $N=4$, $U=0$, $J_{1}=1.0$ and $J_{2}=2.0$.
@@
The Otto cycle operating in the engine mode, by definition, relies on thermal energy for work output; the latter thus vanishes in the limit of zero temperature of the heat baths. However, quantum statistical properties depend crucially on the nature of particles at low temperatures, and this has recently been exploited to design and experimentally realize the Pauli engine [Koch2023]. In this variant of the regular Otto engine, the thermalization strokes are replaced by the so-called Pauli strokes, in which the system undergoes a bosonization or fermionization process. Specifically, rather than thermalizing with a heat bath, the statistical nature of the system is altered by tuning it from a molecular Bose condensate to a Fermi gas or vice versa. By virtue of the Pauli exclusion principle, the system has a higher energy ground state in the fermionic state as compared to the bosonic state. The Pauli engine operates by extracting this difference in ground state energy, defined as the Pauli energy, into useful work. Importantly, the Pauli engine has no classical analogue and is purely driven by quantum statistical phenomena. To this end, we note that the relation between the performance of QTMs and the statistical properties of the system has been explored in a number of other works, especially with respect to fermionic and bosonic statistics [myers20, watanabe20, souza22, myers22, ghosh23, Boubakour2023, Menon2025].
@@
In parallel, it has also been shown in recent times that exotic quantum particles such as anyons [Kundu99, anyonRev08, sternRev08, Iqbal24, Zhang25, Himanshu25, Ghosh25, Yutushui25], including impurity excitations in fractional quantum Hall liquids which exhibit fractional angular momentum and effective anyonic statistics [grass20, baldelli21], and paraparticles [hazzard25, Mostaan25], satisfy statistical properties different from bosons or fermions. The performance of QTMs based on systems satisfying such non-trivial statistics has largely remained unexplored, except recent models based on few-body anyonic systems [Myers2021, manianyonotto24, manianyoncarnot25, Dunlop25]. In this work, we establish how anyonic statistics leads to non-trivial performance enhancement in a many-body QTM which, importantly, can be realised using current experimental platforms. Specifically, we define and analyze a hybrid Anyon-Otto (HAO) cycle (see Fig.1), based on the anyon Hubbard model (AHM) [keilmann11, Longhi12, anyonhubbard15, Tang15, ArcilaForero16, Zhang17, Zhang20, Bartolomei20, Bonkhoff21, Bonkhoff23, greinerExpt24, Dhar2025, Wang25, Bonkhoff2025, Theel25], in which the statistical properties of the particles can be continuously tuned from boson-like to pseudo-fermion-like (satisfy anti-commutation relations only when particles are on different sites). We contrast the operation of the HAO cycle with the Otto cycle in the same temperature and parameter regime except that the statistical properties of the system remain unchanged throughout the Otto cycle. We find that the HAO cycle can produce a finite work output even in the limit of vanishing temperatures of the baths, due to the presence of an anyon energy, defined analogously to the Pauli energy. In the absence of explicit interactions between the anyons, we demonstrate that the HAO cycle can operate in an inverse accelerator mode for finite but small temperatures. This mode, characterized by a heat transfer from colder to hotter bath and a net work output, exhibits an apparent violation of the second law which is reconciled when the work production in the anyonization stroke is appropriately accounted for. In presence of weak interactions between the anyons, we observe that anyonic statistics boosts the performance of the QTM at low temperatures, manifested in the form of higher work output as compared to bosonic or pseudo-fermionic statistics. Finally, we propose how the HAO cycle can be realized experimentally through a feasible extension of the experimental protocol used for the recent realization of the AHM [keilmann11, greinerExpt24, Dhar2025], thus rendering the possibility of direct experimental verification of our results.
II Results
@@
The 1D AHM is realized by introducing a synthetic gauge field in the hopping term of the Bose Hubbard model [anyonhubbard15],
$H=-J\sum_{j=1}^{L}\left(b_{j}^{\dagger}e^{-in_{j}\theta}b_{j+1}+h.c.\right)+\frac{U}{2}\sum_{j=1}^{L}n_{j}\left(n_{j}-1\right),$ (1)
where $L$ is the total number of sites, $b_{j}$ are the bosonic annihilation operators, $n_{j}=b_{j}^{\dagger}b_{j}$ is the number operator, $J$ is the tunneling amplitude, $U$ is the interaction strength, and $\theta$ is a phase parameter. Defining the annihilation operators $a_{j}=e^{-i\theta\sum_{1\leq l\leq j-1}n_{j}}b_{j}$, the equivalent anyon model is obtained,
$H=-J\sum_{j=1}^{L}\left(a_{j}^{\dagger}a_{j+1}+h.c.\right)+\frac{U}{2}\sum_{j=1}^{L}n_{j}\left(n_{j}-1\right),$ (2)
@@
The HAO cycle that we propose here is a four-stroke cycle that inherits operational characteristics from both the Pauli and Otto cycles. Thus, a hallmark of this cycle is its ability to function both as an anyon engine (defined analogously to the Pauli engine) at low temperature and as the standard Otto engine at high temperature. Before defining the HAO cycle, we recall that in the Otto cycle, work and heat are commonly associated with the energy exchange in strokes undergoing exclusively unitary or dissipative dynamics, respectively,
$W_{otto}=\int{\rm Tr}\left[\rho\frac{\partial H}{\partial t}\right]dt,$ (3a)
$Q_{otto}=\int{\rm Tr}\left[\frac{\partial\rho}{\partial t}H\right]dt,$ (3b)
where $\rho$ is the density operator representing the state of the system. In other words, a unitary evolution with a time-dependent Hamiltonian only contributes to the work, while a dissipative evolution, such as thermalization with a heat bath, only results in heat transfer in the Otto cycle.
We now define the HAO cycle as follows. At the start of the cycle, we assume that the system with Hamiltonian $H(\theta_{1},\lambda_{1})$ is in equilibrium with a thermal bath with temperature $T_{A}$ (inverse temperature $\beta_{A}$), and $\lambda$ representing either of the parameters $J$ or $U$ appearing in the Hamiltonian. The initial state is therefore a thermal Gibbs state $\rho_{1}\equiv e^{-\beta_{A}H(\theta_{1},\lambda_{1})}/{\rm Tr}\left[{e^{-\beta_{A}H(\theta_{1},\lambda_{1})}}\right]$. The HAO cycle then consists of the following sequential strokes (see Fig. 1):
@@
    Thermalization B: ($\theta_{1}\to\theta_{2}$, contact with heat bath) - This stroke consists of two sub-strokes – the phase parameter is tuned from $\theta_{1}$ to $\theta_{2}$ followed by thermalization with a heat bath of temperature $T_{B}$. Note that the two sub-strokes can be carried out simultaneously, provided that the phase parameter is changed over a time interval much shorter than the thermalization time-scale. The steady state reached is thus given by, $\rho_{3}=e^{-\beta_{B}H(\theta_{2},\lambda_{2})}/{\rm Tr}\left[{e^{-\beta_{B}H(\theta_{2},\lambda_{2})}}\right]$. The change in the energy expectation value in this stroke has two contributions, one from the energy dissipated to the heat bath and the other from the work done in changing the phase parameter. However, for reasons that we shall clarify below, we associate the total energy change in this stroke with heat transfer,
@@
    Thermalization A: ($\theta_{2}\to\theta_{1}$, contact with heat bath) - In the final stroke, the phase parameter is restored to its initial value $\theta_{1}$ and the subsequent thermalization with heat bath at temperature $T_{A}$ results in the system returning back to its initial state $\rho_{1}$. Thus,
@@
The total work $W=W_{12}+W_{21}$, $Q_{A}$ and $Q_{B}$, satisfies energy conservation (first law) $W=Q_{A}+Q_{B}$. The difference of the HAO cycle defined above with the standard quantum Otto cycle lies in the two modified thermalization strokes. In the latter case, the statistical properties of the quantum model remain unchanged, and thus the heat change is solely associated with the energy gained or lost to the heat baths. In contrast, thermalization in the HAO cycle is accompanied by a explicit change of statistical properties through the parameter $\theta$. It is important to note that ramping the phase parameter modifies the Hamiltonian itself and thus technically amounts to performing a certain amount of work. However, in the anyon equivalent of the model defined in Eq. (2), $\theta$ implicitly controls the statistical properties and hence the equilibrium configuration of the system. Since the purpose of the HAO engine is to convert both thermal energy and anyon energy into useful work, we associate the total energy change in the modified thermalization strokes only with heat energy.
@@
We first consider the operation of the HAO cycle in the non-interacting limit $U=0$ with $\lambda\equiv J$. In this limit, the Hamiltonians at different instants of time during the work strokes commute with each other, implying adiabatic evolution and thus rendering the cycle operation independent of the work stroke duration $\tau$. We also assume that the total number $N$ of particles remain conserved throughout the cycle. In Fig. 2-2, we examine the possible modes of operation for the standard Otto cycle ($\theta_{1}=\theta_{2}=0$), choosing without loss of generality $J_{1}=2.0$, $J_{2}=1.0$, $L=8$ and $N=4$. Depending on the sign of $W$, $Q_{A}$ and $Q_{B}$, and the relative magnitude of $T_{A}$ and $T_{B}$, three modes of cycle operation can be identified - namely the engine mode (E) with $Q_{A},W>0$, $Q_{B}<0$; the refrigerator mode (R) with $Q_{A},W<0$, $Q_{B}>0$ for $T_{A}>T_{B}$, and the accelerator mode (A) with $Q_{A},W>0$, $Q_{B}<0$ for $T_{A}<T_{B}$. The E-R transition can be identified by the line $W=Q_{A}=Q_{B}=0$. For $T_{A},T_{B}\to 0$, the work output and the heat vanish identically.
Having established the modes of operation in the Otto cycle, we now consider the situation $\theta_{1}=\pi$, $\theta_{2}=0$, that is, the statistical properties of the system are altered between bosonic and pseudo-fermionic during the thermalization strokes. From Fig. 2-2, it is evident that in the large temperature limit of the baths, $T_{A},T_{B}\gg 1$, the HAO cycle is identical to the Otto cycle owing to the fact that the anyonic statistics can be well-approximated by Boltzmann statistics in this limit. However, at low temperatures, $T_{A},T_{B}\to 0$, the excess energy resulting from the development of anyonic exclusions begins to manifest itself in the form of the non-zero work output and heat exchanges. This leads to the emergence of an inverse accelerator (IA) mode - heat flows from the colder bath to hotter bath with a net work extraction, $Q_{A},W>0$, $Q_{B}<0$ for $T_{A}<T_{B}$. As we shall show later, the IA mode does not violate the second law of thermodynamics since the anyon energy gained by the system during the anyonization process involves an additional work cost, which when included in the definition of $W$ restores the second law.
The finite work output in the limit of vanishing temperature manifests itself not only in the pseudo-fermionic limit $\theta_{1}=\pi$, but for any $\theta_{1}\neq 0$. This is demonstrated in Fig. 3, where we observe that the average work output per particle $W/N$ increases monotonically with increasing $\theta_{1}$ in the inverse accelerator mode ($T_{A}=0.2$, $T_{B}=0.4$), provided $T_{A}\to 0$. Similarly, the work input decreases with increasing $\theta_{1}$ in the accelerator mode ($T_{A}=0.2$, $T_{B}=0.8$) in the same limit of $T_{A}$. Note that it is sufficient for only $T_{A}$ to be small as the anyonization process occurs in contact with the heat bath of temperature $T_{A}$.
@@
We now consider the situation in which work is performed by ramping the interaction parameter $U$ during the work strokes, holding $J$ to a constant value. In this case, the Hamiltonian does not commute with itself at different times during the work strokes, which leads to non-adiabatic excitations for finite stroke duration $\tau$. However, for simplicity, we shall assume that $\tau\to\infty$ as the results discussed in the following do not change qualitatively in the presence of non-adiabatic excitations. In order to ensure that the anyon energy is not affected by the interaction, we choose $U_{1}=0$. The work output $W$ at low temperature, $T_{A}=T_{B}=0.1$, as a function of the statistical parameter $\theta_{1}$ is shown in Fig. 4 for different interaction strengths $U_{2}$. In the weakly interacting limit $U\ll J$, we observe that $W$ is maximized for $\theta_{1}=\theta^{*}$ such that $0<\theta^{*}<\pi$, for $N\geq L/2$. However, as the interaction strength is increased, the maxima shifts towards the pseudo-fermionic limit $\theta^{*}=\pi$.
@@
The higher probabilities of double or greater occupancy in the bosonic and pseudo-fermionic limit can be understood by reverting back to the bosonic Hamiltonian defined in Eq. (1). Since the hopping amplitude acquires a density dependent phase for any $\theta\neq 0$, the system minimizes energy by preferring lesser contributions from configurations in which a given site is doubly or more occupied. For less than half-filling $N<L/2$, this is easily achieved as there are sufficient number of empty sites to satisfy lesser or equal to unit occupancy per site. Thus, the probability for higher occupancy decreases monotonically with increasing value of $\theta$. For $N\geq L/2$, the situation is not straightforward and the system is forced to choose configurations with greater than unit occupancy for some sites. However, as $\theta\to\pi$, the density dependence phase reduces to unity even for higher occupancy, and thus the probability of more than unit occupancy increases again. To support the above argument, we plot in Fig. 5 the probability $P(n_{j})$ that the $j^{th}$ site is occupied by $n_{j}$ particles for a couple of nearby sites $j=3,4$ in the system with $L=8$, $J=1.0$ and $U=0$. Although $P(n_{j}=0,1)>P(n_{j}=2,3,4)$ for both the sites, $E_{per}$ has no contributions from the zero or unit occupancies as the interaction energy vanishes identically for $\braket{\psi_{0}|n_{j}|\psi_{0}}=0,1$. It can be seen from Fig. 5 that the $P(n_{j}=2)$ decreases initially as $\theta$ is increased from zero, before increasing again on approaching $\theta=\pi$. Importantly, the increase in the value of $P(n_{4}=2)$ is accompanied by an increase in the value of $P(n_{3}=0)$ and a decrease in the value of $P(n_{3}=1)$, suggesting that double occupancy is preferred at the expense of unit occupancy in the limit $\theta\to\pi$.
Reconciliation with the second law
In the previous sections, we have argued that the apparent violation of the second law of thermodynamics in the IA mode can be reconciled if the work associated with the buildup of anyon energy is accounted for. To show this explicitly, we modify the definition of work in Eq. (4) and Eq. (6) to incorporate the energy change associated with the anyonization and bosonization processes. Likewise, we subtract the same from the heat exchanges to ensure consistency with the first law. Furthermore, in this case, we assume that the anyonization/bosonization process is carried out much faster than the time-scale of thermalization with the baths. This ensures that energy change during the anyonization/bosonization process is only due to work performed by explicit change of the parameter $\theta$ in the Hamiltonian. Thus, the redefined work and heat are, $\overline{W}_{12(21)}={W}_{12(21)}-\Delta_{B(A)}$, $\overline{W}=\overline{W}_{12}+\overline{W}_{21}$, $\overline{Q}_{A(B)}=Q_{A(B)}-\Delta_{A(B)}$, where